### v1

# Making Large Language Models Better Knowledge Miners for Online Marketing with Progressive Prompting Augmentation

Abstract.
Nowadays, the rapid development of mobile economy has promoted the flourishing of online marketing campaigns, whose success greatly hinges on the efficient matching between user preferences and desired marketing campaigns where a well-established Marketing-oriented Knowledge Graph (dubbed as MoKG) could serve as the critical “bridge” for preference propagation. In this paper, we seek to carefully prompt a Large Language Model (LLM) with domain-level knowledge as a better marketing-oriented knowledge miner for marketing-oriented knowledge graph construction, which is however non-trivial, suffering from several inevitable issues in real-world marketing scenarios, i.e., uncontrollable relation generation of LLMs, insufficient prompting ability of a single prompt, the unaffordable deployment cost of LLMs. To this end, we propose PAIR, a novel Progressive prompting Augmented mIning fRamework for harvesting marketing-oriented knowledge graph with LLMs. In particular, we reduce the pure relation generation to an LLM based adaptive relation filtering process through the knowledge-empowered prompting technique. Next, we steer LLMs for entity expansion with progressive prompting augmentation, followed by a reliable aggregation with comprehensive consideration of both self-consistency and semantic relatedness. In terms of online serving, we specialize in a small and white-box PAIR (i.e.,LightPAIR), which is fine-tuned with a high-quality corpus provided by a strong teacher-LLM. Extensive experiments and practical applications in audience targeting verify the effectiveness of the proposed (Light)PAIR  [^1] Source code and datasets will be released upon acceptance..
Knowledge Graph; Online Marketing Campaign; Large Language Model
* Equal contribution.
${}^{\dagger}$ Corresponding author.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Artificial intelligence
1. Introduction
Nowadays, the rapid growth of mobile economy has established the prosperity of online marketing campaigns (Yu et al., 2021; Chen et al., 2021; Cai et al., 2023; Liu et al., 2023), which attracts a variety of Internet companies or merchants to promote their products / services. Taking Alipay as an example, as an integrated service platform in China, numerous merchants conduct marketing campaigns (e.g., offering coupons, vouchers, and bonuses) through so-called mini-apps each day (Zhuang et al., 2020; Zhu et al., 2021), in which efficient information delivery plays a crucially important role. Fundamentally, the core is to perform intelligent matching between user preference and desired campaigns(deWet and Ou, 2019; Liu et al., 2020; Rahman et al., 2023; Yan et al., 2023), where a well-established Marketing-oriented Knowledge Graph (dubbed as MoKG) could serve as the critical “bridge”, allowing flexible and effective preference propagation. Although SupKG (Zang et al., 2023) [^2] SupKG involves more than ten millions of entities, more than one hundred relations, and more than one hundred million of triplets. has made considerable efforts in connecting users and various services in Alipay, product-related entities with hierarchical and spatiotemporal relations are at its heart, while MoKG emphasizes more on the marketing-related relations between users and merchants (A snapshot of MoKG is illustrated in Fig. 1), as a kind of complementary knowledge to enrich the original SupKG graph (More details in Sec. 4.3.3).
Intuitively, we could directly graft the construction pipeline of SupKG, which is a mixture of multiple text mining techniques (e.g., NER (Yadav and Bethard, 2018; Rijhwani et al., 2020; Mueller et al., 2020; Das et al., 2022; Zhang et al., 2023a) and RE (Zhu et al., 2019; Chen et al., 2023; Gururaja et al., 2023; Xu et al., 2023b)), to supplement marketing-related entities and relations. Unfortunately, its success greatly hinges on manual labeling and evaluation, which is rather expensive. On the other hand, Large Language Models (LLMs) (e.g., ChatGPT (OpenAI, 2023a), GPT4 (OpenAI, 2023b), Llama  (Touvron et al., 2023a, b), ChatGLM (Du et al., 2022; Zeng et al., 2023)) have shown remarkable and even superhuman capabilities on various NLP tasks. Learning from a vast corpus of Internet text, LLMs have encoded a huge amount of open-world knowledge (Petroni et al., 2019; AlKhamissi and Ghazvininejad, 2022; Cohen et al., 2023), which are naturally excellent helpers to harvest factual information for knowledge graphs (Hao et al., 2023; Cohen et al., 2023; Zhu et al., 2023). However, the effectiveness of LLM based mining approaches cannot be always guaranteed in real-world marketing scenarios, since pure LLMs may be unfamiliar with the marketing-oriented entities and relations(Sun et al., 2023; Dhuliawala et al., 2023). Thus, can we bridge both worlds, i.e., carefully prompting LLMs with prior knowledge (e.g., SupKG or other knowledge) as a better marketing-oriented knowledge miner ?
Figure 1. An example of the marketing-oriented knowledge graph.
With this in mind, we decompose the whole mining process of MoKG as the following three stages: Prior Knowledge $\rightarrow$ Relation Generation $\rightarrow$ Entity Expansion. With the help of prompting engineering, domain-level prior knowledge could be easily injected into LLMs for relation generation and entity expansion, whereas the performance is still far from optimal for real-world marketing-oriented scenarios due to the following unresolved issues: (i) Pure relation generation is uncontrollable. Despite the impressive performance of LLMs in various generative tasks, it is undeniable that the capabilities of LLMs can get out of control very quickly in relation exploration. In other words, LLMs may inevitably generate several useless / fallacious (Zhang et al., 2023b, c; Jiang et al., 2023) relations for given entities (e.g., Take the entity “The Three Body” in Fig. 1 as an example, the LLM would tell several undesired relations towards marketing, e.g., “Related scientists” and “Related organizations”.) so that the following entity expansion would be greatly confused. (ii) A single prompt is insufficient. Intuitively, the performance of entity expansion depends highly on the quality of the prompt used to steer the LLMs. Unfortunately, due to the variety of prior knowledge, coupled with the sensitivity of prompting under slight modifications (Shin et al., 2020; Chakraborty et al., 2023; Arora et al., 2023), a single painstakingly crafted prompt may not always produce the desired results. (iii) A white-box and smaller model is preferred. Generally, real-world marketing scenarios involve millions of seed entities, which brings far beyond affordable cost to directly serve (or call open API of) a performant LLM (e.g., GPT 3.5). On the other hand, the growing concerns for privacy demand to circumvent a black-box LLM in realistic applications.
To address the above problems, we propose a novel Progressive prompting Augmented mIning fRamework for harvesting marketing-oriented knowledge with LLMs (dubbed as PAIR). To perform relation generation in a controllable manner, we restrict this procedure in a finite relation space, which is thus reduced to an LLM based relation filtering process (i.e., adaptively determining which relations deserve expansion for each entity). In terms of entity expansion, we steer LLMs with well-designed progressive prompting augmentation (i.e., from coarse-grained to fine-grained prompts), followed by a reliable aggregation with comprehensive consideration of both self-consistency and semantic relatedness. To remedy the shortcoming of LLMs in mastering marketing-oriented knowledge, we collect relevant evidence from SupKG and available knowledge bases to empower LLMs in both relation filtering and entity expansion stages. And then, we specialize in a small and white-box model (dubbed as LightPAIR) with a high-quality corpus provided by a strong teacher-LLM involving 175 billion parameters. At last, extensive experiments and practical applications in audience targeting verify and justify the effectiveness of the proposed (Light)PAIR.
2. Related work
2.1. Traditional knowledge graph construction
Currently, traditional construction process of well-known knowledge bases / KGs usually requires heavy human labor for labeling and evaluation or multiple text mining-based procedures. In the line of manual labor-intensive research (Bollacker et al., 2008; Hoffart et al., 2013), WordNet (Fellbaum, 1998) is a database that links words with semantic relations, ConceptNet (Speer et al., 2017) is a large commonsense knowledge graph consists of a large set of knowledge triples while ATOMIC (Sap et al., 2019) is a crowd-sourced social commonsense KG of inferential knowledge which is organized as typed if-then relations. On the other hand, text mining-based approaches usually consist of sub-procedures such as named entity recognition(Lample et al., 2016; Das et al., 2022; Zhang et al., 2023a), relation extraction(Zhu et al., 2019; Chen et al., 2023; Gururaja et al., 2023), semantic parsing(Kamath and Das, 2019; Li et al., 2022; Zhang et al., 2023d) that require to collect different text corpus to train independent model for each procedure. Among them, Quasimodo (Romero et al., 2019), TransOMCS (Zhang et al., 2020), DISCOS (Fang et al., 2021), ASCENT (Nguyen et al., 2021), SupKG (Zang et al., 2023) construct commonsense or practice-oriented knowledge graphs by making use of linguistic patterns and huge manual processing.
Open issues. These research works usually hinge on collecting different text samples to train multiple independent models for each procedure (e.g., named entity recognition and relation extraction) and require huge manual resources for labeling. However, the limited manual resources and text corpora usually restrict the scale of constructed knowledge graph.
2.2. Language model-driven knowledge graph construction
For works in this direction, they tend to 1) distill knowledge graphs from LMs, 2) incorporate the power of LMs to enhance knowledge graph construction. In COMET (Bosselut et al., 2019), implicit knowledge from deep pre-trained language models is transferred to generate explicit knowledge in commonsense knowledge graphs. MAMA (Wang et al., 2020) proposes to use learned knowledge stored in pre-trained LMs in an unsupervised manner to construct a knowledge graph and compares the quality of it with human-labeled knowledge graph. BertNet (Hao et al., 2023) generates multiple prompts for KG generation and weights them by the likelihood of prompts given by an LLM. In LMCRAWL (Cohen et al., 2023), relation paraphrases and object paraphrases are utilized for KG generation. On the other hands, (Trajanoska et al., 2023) compares the advanced LLM e.g., ChatGPT with the specialized pre-trained models e.g., REBEL for entity and relation extraction.
Open issues. In these research works, they tend to generate knowledge graphs or knowledge bases towards world knowledge / contain knowledge stored in the parameters of LMs, while in industrial scenarios, for example, commercial marketing scenarios, usually pure LMs may be unfamiliar with the marketing-oriented entities and relations hence those approaches cannot meet the requirements.
3. The Proposed PAIR Framework
In this section, we dive into the elaboration of the proposed PAIR, a novel Prompting Augmented mIning fRamework for harvesting marketing-oriented knowledge with LLMs.
3.1. Formulation
Essentially, a mining procedure of a knowledge graph could be formulated as follows: with the given source entity $s$, the likelihood of target entity $t$ and its corresponding relation $r$ is calculated as
(1) $\begin{split}\mathcal{P}(r,t|s)&=\int_{\kappa}\mathcal{P}(\kappa|s)\mathcal{P}(r,t|s,\kappa)\\ &=\int_{\kappa}\mathcal{P}(\kappa|s)\mathcal{P}(r|s,\kappa)\mathcal{P}(t|s,\kappa,r)\\ \end{split}$
  • $\mathcal{P}(\kappa|s)$: Prior Knowledge. Intuitively, the knowledge miner commonly struggles to perceive the real meaning of the given entities in the marketing-oriented applications, e.g., sneaker-related entities might be preferred with the given entity “Air”. Hence, the usage of prior knowledge $\kappa$ occupies a critical position in the guidance of knowledge mining.
  • $\mathcal{P}(r|s,\kappa)$: Relation Filtering. Based on the enormous knowledge $\kappa$, the pure process of relation generation is ideal but generally uncontrollable, especially in marketing-oriented applications where a bundle of relations might be undesired, e.g., relation “Related scientists” w.r.t. entity “The Three Body”. Therefore, we resort to the powerful capability of LLMs for filtering desired relations $r$ from the pre-defined relation set $\mathcal{R}$.
  • $\mathcal{P}(t|s,\kappa,r)$: Entity Expansion. With careful prompting engineering, we communicate with LLMs to complete the entity expansion task. However, due to the variety of prior knowledge $\kappa$, coupled with the sensitivity of prompting under slight modifications, a single painstakingly crafted prompt may not always produce the desired results. Instead, we prompt the LLMs for entity expansion with multiple prompting augmentations via a reliable aggregation strategy.
In the following, we will delve into the instantiation of each part with the help of the recent emerging LLMs.
Figure 2. Overall architecture of PAIR.
3.2. The Usage of Prior Knowledge
In practice, LLMs commonly struggle to perceive the real meanings of the given entities due to the lack of domain-level knowledge, which severely hinders the capabilities of LLMs in both relation filtering and entity expansion stages. Hence, it is of crucial importance to introduce prior knowledge to specialize in a general LLM for marketing-oriented scenarios. In short, for each entity $e$, we mainly collect the following two complementary knowledge to serve as key evidence of LLMs for reasoning.
  • Structural Knowledge. The collection of structural knowledge benefits from the establishment of SupKG (Zang et al., 2023), which involves abundant side information of entities, including brand, category and so on. In order to maintain the reliability of structural knowledge and introduce as less noise as possible, for each entity, we only collect the one-hop neighbors in the SupKG, coupled with the entity type for disambiguation (e.g., The type of entity “Apple” could be “Brand” or “Food”).
  • Descriptive Knowledge. Due to the incomplete and noisy fact of SupKG, we are unable to collect all the convincing evidence for steering LLMs. Hence, we resort to descriptive knowledge from public (i.e., Wikipedia) and private knowledge bases. Taking the entity “The Three Body” as an example, the SupKG only provides its entity type “Book” and its category “Science Fiction”, while the descriptive summary in Wikipedia offers stronger evidence (i.e., author information and brief introduction) to support effective knowledge mining.
3.3. LLM as a Relation Filter
As mentioned above, to make the whole procedure of relation generation under control, we restrict the process to a finite relation space. Hence a relation filter is encouraged since not all relations in the whole relation set $\mathcal{R}$ are helpful / appropriate for a given entity to expand marketing-oriented entities (e.g., given the brand entity “Apple”, the relation “Related Food” may confuse the LLMs for entity expansion). Therefore, we resort to the powerful capability of LLMs for filtering desired relations from the pre-defined relation set $\mathcal{R}$. Given a source entity $s$ with an LLM $\mathcal{M}$, we define the relation filtering as follows:
(2) $s\xrightarrow[(1)]{\phi(\cdot;\mathcal{K}),\mathcal{R}}\mathcal{R}^{\mathcal{T}}_{s}\xrightarrow[(2)]{\mathcal{M}(\cdot;\rho^{\mathcal{R}}(s,\kappa))}\mathcal{R}^{\mathcal{F}}_{s}.$
  1. (1)
[…]
  2. (2)
    Filtering. We come up with applying the LLM as a relation filter with structural / descriptive knowledge augmented prompting engineering $\rho^{\mathcal{R}}(s,\kappa)$, i.e., denoted as “{Struc_KG}” and “{Desc_KG}” in Fig. 2. With the designed prompt, an LLM would provide i) desired relations $\mathcal{R}_{s}^{\mathcal{F}}$ that align with the source entity and corresponding prior knowledge, allowing for more reasonable entity expansion; ii) a potential target entity corresponding to each desired relation, which could be regarded as a kind of inheritable knowledge for entity expansion.
3.4. LLM as an Entity Expansionist
Due to the variety of prior knowledge, coupled with the sensitivity of prompting under slight modifications, simply employing a single painstakingly crafted prompt may not always produce the desired results. To this end, we steer LLMs with careful progressive prompting augmentation, followed by a reliable aggregation that takes both self-consistency and semantic relatedness into consideration. Given a source entity $s$ and the corresponding relation $r$, a pipeline of the entity expansion with an LLM $\mathcal{M}(\cdot)$ is formulated as follows:
(3) $(s,r)\ \ \ \begin{matrix}\\ \xrightarrow[]{\mathcal{M}(\cdot;\rho^{\mathcal{E}}(s,\kappa_{1},r))}\mathcal{T}^{1}_{s,r}\\ \xrightarrow[]{\mathcal{M}(\cdot;\rho^{\mathcal{E}}(s,\kappa_{2},r))}\mathcal{T}^{2}_{s,r}\\ \xrightarrow[(1)]{\mathcal{M}(\cdot;\rho^{\mathcal{E}}(s,\kappa_{...},r))}\mathcal{T}^{...}_{s,r}\end{matrix}\xrightarrow[(2)]{\mathcal{A}(;)}\mathcal{T}^{\mathcal{F}}_{s,r}.$
(1) Progressive prompting augmentation. It is designed for navigating an LLM $\mathcal{M}$ to different aspects of prior knowledge $\mathcal{\kappa}$, making the expansion procedure more diverse and robust. In particular, we mainly prepare three kinds of knowledge: structural knowledge $\kappa^{\mathcal{S}}$, descriptive knowledge $\kappa^{\mathcal{D}}$ and inheritable knowledge from relation filtering $\kappa^{\mathcal{I}}$, i.e., denoted as “{Struc_KG}”, “{Desc_KG}” and “{Inher_KG}” in Fig. 2. By incorporating various knowledge, we progressively construct prompts from coarse-grained to fine-grained manner as $\{\rho^{\mathcal{E}}(s,r;\kappa^{\mathcal{S}}),\rho^{\mathcal{E}}(s,r;\kappa^{\mathcal{D}}),\rho^{\mathcal{E}}(s,r;\kappa^{\mathcal{I}}),\rho^{\mathcal{E}}(s,r;\kappa^{\mathcal{S}},\kappa^{\mathcal{D}}),\cdots,\\ \rho^{\mathcal{E}}(s,r;\kappa^{\mathcal{S}},\kappa^{\mathcal{D}},\kappa^{\mathcal{I}})\}$, and query a performant LLM repeatedly for stability (Fig. 2 presents one of the prompts for entity expansion, while more detailed examples of prompting augmentation will be shown in Sec. A.). As a result, we define the mined entities from the progressive prompt augmentation as $\left\{\mathcal{T}^{1}_{s,r},\mathcal{T}^{2}_{s,r},\cdots,\mathcal{T}^{...}_{s,r}\right\}$.
(2) Reliable aggregation. After entity expansion with progressive prompting augmentation, we aim at seeking an aggregation mechanism $\mathcal{A}(\cdot)$ in a reliable manner to obtain the final entity expansion result, which takes into consideration the following main aspects:
[…]
  • Self-consistency of the knowledge. Specifically, it aims at finding the most consistent entities when the LLM performs reasoning in multiple ways. Intuitively, when an entity appears across multiple expansion results, it indicates the higher confidence of LLMs towards the particular target entity, which can then be utilized to reward the highly self-consistent target entities. Here, we denote the total number that $t$ appears across multiple entity expansion procedure as $\sum_{k}\mathbb{I}(t\in\mathcal{T}_{s,r}^{k})$, where $k$ indicates a specific round in progressive prompt augmentation and $\mathbb{I}(t\in\mathcal{T}_{s,r}^{k})$ returns $1$ if $t$ in $\mathcal{T}_{s,r}^{k}$ else 0.
By integrating both aspects, we define the final score for aggregation as follows:
(4) $\tau_{s,r,t}=\log(1+\sum_{k}\mathbb{I}(t\in\mathcal{T}_{s,r}^{k}))\cdot\text{MLP}(\bm{x}_{s,r,t})$
Finally, given a source entity $s$ and relation $r$, we only keep top $K$ [^4] In our experiment, we set $K$ to 8 w.r.t. the trade-off between multiple metrics. mined entities with the highest scores from aggregation as the entity expansion result, denoted as $\mathcal{T}^{\mathcal{F}}_{s,r}$.
Figure 3. The finetuning and inference pipeline of LightPAIR.
3.5. LightPAIR: Towards Deployment with a Strong Teacher-LLM
In general, the effectiveness of PAIR greatly hinges on the remarkable advancements of LLMs, whereas, equipping PAIR with today’s state-of-the-art LLMs (i.e., GPT 3.5 with 175 billion parameters) is rather difficult even impractical to most of the real-world applications due to far beyond the affordable cost of memory and compute. Meanwhile, inspired by the common insight that the corpus generated by a strong teacher-LLM is highly valuable for enhancing smaller task-specific models via careful fine-tuning(Hsieh et al., 2023; Shi et al., 2023; Xu et al., 2023a), we seek to conduct a lightweight and white-box pipeline of PAIR (i.e.,LightPAIR) for large-scale MoKG mining from the complete PAIR. In particular, we collect a group of seed entities (i.e., around 4K $\sim$ 10K) from SupKG and feed them into the complete PAIR. Subsequently, we perform a human evaluation (detailed in Sec.4.1) on the output entity set, followed by bad-case filtering, e.g., remove relations / target entities that are annotated as incorrect by human annotators. To obtain a high-quality corpus for fine-tuning the smaller language models, we follow the prompt template of relation filtering and entity expansion to construct questions, coupled with corresponding responses. Based on the MoKG-oriented corpus, current lightweight LLMs (e.g., Bloomz (Muennighoff et al., 2022), GLM (Du et al., 2022), ChatGLM (Zeng et al., 2023), Baichuan-7B (Baichuan, 2023)) could be easily fine-tuned (i.e., supervised fine-tuning, short as SFT) via the parameter-efficient Lora (Hu et al., 2022) strategy. Finally, we inject the smaller fine-tuned LLM into the relation filtering and entity expansion procedure of PAIR, and adopt the LightPAIR for mining large-scale MoKG.
3.6. Discussion: Diving into the Progressive Prompting Augmentation
Merits From the specific examples of progressive prompting augmentation presented in Sec. A, we can find that when compared with the non-augmented prompt output, the progressive prompting augmentation has the following virtues:
  • Triggering a more reliable knowledge mining in the industrial setting. Usually, pure LLMs may be unfamiliar with many marketing-oriented entities and relations such that they tend to hallucinate in their output, e.g., ”Uncle Fruit power bank” in the case of entity ”Uncle Fruit”, however, the appropriately injected knowledge can provide the right context for LLMs and steer them to the right reasoning direction, hence the output can be closely related to a brand that sells fruit products.
  • Capable of navigating an LLM to different aspects. Take ”Uncle Fruit” as an example, since different knowledge has its emphasis, the output of an LLM w.r.t. different prior knowledge varies from each other (e.g., the output of prompt 1 focuses on common fruit products while the output of prompt 2 focuses more on the product of service) and results in a more diversified output, which we carefully evaluate with two metrics in a coarse-grained to a fine-grained manner and encouraging results can be found in the “Experiments” part.
About time consumption Intuitively, the progressive prompting augmentation requires multiple calls to LLMs, incurring more time consumption compared to the original methods. However, this additional time investment is deemed acceptable due to the following three reasons: i) The progressive prompting augmentation earns remarkable performance gains compared to original prompting methods, detailed in Table 1. ii) The knowledge mining task solely relies on offline inference, and there is no requirement for daily inference. As a result, the increased time consumption has no impact on downstream applications. iii) Since LightPAIR is capable of being deployed on a single A100, we can seamlessly accomplish the overall inference task by utilizing multiple GPUs, thereby meeting the time requirements efficiently.
Table 1. Overall performance comparison of different methods towards mining MoKG. We mark the best and runner-up performance in bold and “-” means these methods fail to mine novel entities.
MoKG-181 MoKG-500
Methods Accuracy Novelty Diversity Accuracy Novelty Diversity
AEE ILAD AEE ILAD
KG Completion BERT 58.4% - 43.0 4.86 57.7% - 42.7 4.97
TRMP 91.1% - 13.8 3.12 91.3% - 14.1 3.07
KG Construction LMCRAWL 86.3% 41.2% 36.3 5.98 85.2% 41.7% 37.1 6.00
COMET 86.7% 35.9% 26.1 5.06 85.9% 34.6% 25.3 4.95
OURS PAIR 90.1% 40.4% 43.7 5.88 90.7% 43.6% 42.8 5.78
 - Agg 88.7% 39.6% 30.8 5.89 88.9% 36.4% 31.3 5.84
 - Agg & Pr 86.9% 36.8% 30.8 5.82 87.2% 34.2% 31.4 5.87
 - Agg & Pr & Rf 84.9% 39.2% 46.3 6.15 84.3% 39.4% 47.2 6.20
4. Experiments
4.1. Experimental Setup
4.1.1. Datasets
Since PAIR aims to mine diverse marketing-oriented knowledge using a group of seed entities and a candidate relation set, we randomly collect two seed entity sets from the well-established SupKG. These entity sets consist of 181 entities (i.e.,MoKG-181) and 500 entities (i.e.,MoKG-500). With domain knowledge, we meticulously gather 105 relations (e.g., “Related Food”, “Issue Voucher”, “Prize of Food” and “Brand of Product”) as the candidate relation set for entity expansion.
4.1.2. Baselines and Variants
We mainly compare PAIR with the following baselines and variants:
  • KG Completion based, which attempts to mine knowledge in the established SupKG through textual matching and structural matching: i) BERT (Devlin et al., 2019), which completes an existing knowledge graph by textual matching using semantic-based representation. ii) TRMP (Yang et al., 2023), which constructs a knowledge graph by retrieval and ranking of entity pairs in a progressive manner and ensembles the entity representation across time to enhance the knowledge graph stability.
  • KG Construction based, which aims at exploring commonsense or open-world knowledge with deep pre-trained language models: i) COMET (Bosselut et al., 2019), which learns from existing knowledge bases and generates novel nodes and edges in commonsense descriptive natural languages. ii) LMCRAWL (Cohen et al., 2023), which designs multiple prompts to mine a knowledge graph out of a large language model (GPT 3.5 in this paper) by following the procedure of Subject Paraphrasing $\rightarrow$ Relation Generation $\rightarrow$ Relation Paraphrasing $\rightarrow$ DK Object Generation.
  • Variants of PAIR: To further examine the merits of each well-designed part in PAIR, we prepare three variants: i) PAIR -Agg, which discards the Aggregation operation and it is equivalent to the most performant version of Progressive prompting. ii) PAIR -Agg & Pr, which discards the Aggregation operation and the Progressive prompting and it is equivalent to directly utilize the output of Relation Filtering. iii) PAIR -Agg & Pr & Rf, which discards the Aggregation operation, the Progressive prompting and the Relation filtering and it is equivalent to directly utilize an LLM for the task of knowledge mining.
For PAIR, we equip it with today’s state-of-the-art LLM (i.e., GPT 3.5 with 175 billion parameters). During the progressive prompting augmentation process, we call the LLM for each specific prompt for 3 times. For reliable aggregation, the KG-BERT (Yao et al., 2019) consists of 110 million parameters, which is the same as BERT (Devlin et al., 2019).
4.1.3. Evaluation Protocols and Metrics
We mainly evaluate the quality of the knowledge mining with human annotation, where three annotators are involved using a correctness / incorrectness judge for each extracted knowledge tuple. A tuple is considered “accepted” if at least two annotators deem it to be true knowledge, and “rejected” if at least two annotators vote it as false. To ensure the credibility of human evaluation, we initially gather all predicted tuples from both baselines and our method. These tuples are then randomly shuffled before being presented for human annotation. In this way, the annotators are unaware of which model a specific knowledge tuple is derived from, which helps to ensure the fairness in the evaluation process.
Overall, we comprehensively measure the quality of mined knowledge from the following aspects: i) Accuracy refers to the proportion of accepted tuples. ii) Novelty refers to the proportion of extracted entities that do not appear in the SupKG. iii) Diversity is evaluated with two metrics from coarse-grained to fine-grained manner. AEE (Average Entity Expansion) refers to the average number of target entities expanded from seed entities. ILAD (Intra-List Average Distance) (Wu et al., 2022) is a distance-based diversity metric, calculating the average pair-wise (Euclidean) distance among all the target entities given a source entity in representation space [^5] The representations are derived from BERT that are pre-trained on Wikipedia and SupKG.. In general, when evaluating the quality of mined knowledge, the accuracy metric comes first. With accuracy guarantee, we hope the mined knowledge with higher novelty and diversity thus better facilitating downstream tasks in real-world environment.
4.2. Performance Overview
In this part, we report the evaluation results of different knowledge mining approaches for mining MoKG in Table 1 and the major findings can be summarized as follows:
  • (Superior performance) The proposed PAIR consistently achieves remarkable performance across all datasets and key metrics, indicating its effectiveness in mining marketing-oriented knowledge graphs. Besides, beyond its considerable performance in accuracy metric, its superior performance in novelty and different aspects of diversity further indicates that the harvested knowledge graph can provide considerable complementary knowledge to the original knowledge graph, i.e., SupKG, thus further facilitating broader applications to downstream tasks in real-world environments.
  • (Compared with KG completion based methods) It is worthwhile to note that KG completion based methods, especially TRMP, yield impressive performance w.r.t. the accuracy metric. However, these methods fail to offer extra knowledge since they fail to mine novel entities, which is unacceptable for the mining of MoKG in real-world applications.
  • (Compared with KG construction based methods) Not surprisingly, KG construction based methods (i.e., LMCRAWL and COMET) generally bring a considerable amount of novel knowledge with the help of deep pre-trained language models. However, the proposed PAIR still surpasses them, attributed to the following well-designed components tailored for mining MoKG: adaptive relation filtering and progressive prompting augmentation, empowered by domain-level knowledge (More case studies are presented in Section 4.3.2).
  • (Ablation studies of PAIR) Clearly, the performance of PAIR would drop a lot once the corresponding module is discarded, indicating that all of our elaborate designs occupy critical positions in mining high-quality and high-diversity MoKG for online marketing. It is noted that the variant PAIR -Agg & Pr & Rf (namely basic prompting method which means directly utilizing an LLM) ranks highest w.r.t. the diversity metric, demonstrating the ability of a performant LLM in finding potential target entities for knowledge graph construction. However, directly utilizing an LLM for this task results in a rapid decrease of accuracy and novelty as compared to PAIR and its other variants, which further justifies our model design.
Table 2. Performance evaluation of LightPAIR instantiated by different student LLMs. We mark the best and runner-up performance of student LLMs in bold.
Models PAIR LightPAIR
LLMs(Size) GPT 3.5(175B) GLM(10B) GLM(10B) Bloomz(7B) ChatGLM2(6B) Baichuan2(7B) Qwen2(7B)
SFT ? -
Accuracy 90.1% 79.2% 89.0% 80.8% 86.3% 90.3% 80.6%
Train- Novelty 40.4% 39.9% 31.0% 29.0% 28.8% 31.5% 25.8%
25K Diversity AEE 43.7 27.44 35.71 48.47 39.21 41.08 25.02
ILAD 5.88 6.14 5.77 6.12 5.82 5.96 5.74
Accuracy 90.1% 79.2% 90.7% 82.7% 88.9% 92.2% 80.6%
Train- Novelty 40.4% 39.9% 31.5% 30.2% 29.5% 34.3% 29.1%
100K Diversity AEE 43.7 27.44 37.10 48.03 38.94 41.13 25.89
ILAD 5.88 6.14 5.83 6.14 5.88 5.79 5.49
4.3. Further Probe
4.3.1. LightPAIR with Different Student-LLMs
In this part, we investigate into the effectiveness of LightPAIR, augmented by different student-LLMs.
Data collection for SFT. With the help of GPT 3.5, we generate two datasets for fine-tuning the smaller LLMs towards relation filtering and entity expansion, consisting of 25K (i.e., Train-25K) and 100K (Train-100K) samples constructed from (4K-10K) seed entities , respectively. In detail, we carefully choose various types of entities and relations from SupKG and utilize the teacher LLM (GPT 3.5) to do relation filtering, and entity expansion by progressive prompting and reliable aggregation. Subsequently, we select 3 annotators to perform a thorough manual verification of the results derived from the teacher LLM, which ensures the creation of a high-quality training dataset. On the other hands, we utilize the MoKG-181 as the evaluation dataset in this experiments   [^6] Due to the space limit, we omit the experimental results on MoKG-500 in this paper. However, it is worth mentioning that similar trends are observed as well..
Experimental settings of student-LLMs. With the collected training corpus derived from the strong teacher-LLM, we adopt several currently popular open-source LLMs to instantiate the LightPAIR, i.e., GLM (Du et al., 2022), Bloomz (Muennighoff et al., 2022) and ChatGLM2 (Zeng et al., 2023), Baichuan-7B (Baichuan, 2023), Qwen-7B (Bai et al., 2023), and set the optimizer, maximum context size, batch size, and learning rate to Adam, 4096, 8 and 5e-5 respectively.
Results and analysis. From Table 2, by comprehensively taking multiple metrics into consideration we could find that LightPAIR with GLM(10B) and Bainchuan2(7B) could achieve comparable performance of the complete PAIR with GPT 3.5(175B), implying that high-quality corpus generated by a strong teacher-LLM can greatly enhance smaller task-specific models. Besides, as the training corpus accumulates, an increasing trend among multiple metrics can be observed which also depicts the significance of collecting high-quality task-oriented corpus for enhancing task-specific models. When compared with those student-LLMs with SFT (short for Supervised Fine-Tuning), though a pure student-LLM without SFT ranks highest w.r.t. ILAD, the lowest accuracy of the collected knowledge prohibits its application in real-world industrial environments.
Figure 4. Illustration of the knowledge graph enriched by the PAIR.
Table 3. Case study towards the usage of prior knowledge in PAIR. We bold the hallucinative entities and fallacious entities in red and blue, respectively.
Source entity Relation type Target entities
Mi Xiao Quan related media name w/o knowledge Journey to the West (西游记)
(米小圈上学记) w/ knowledge Tom and Jerry(猫和老鼠), Boonie Bears(熊出没)
CKA target audience w/o knowledge System Administrator (系统管理员)
(空手道协会) w/ knowledge Karate Enthusiasts(空手道迷), Wushu Master(武术师)
Uncle Fruit related brand w/o knowledge Fruit Education(果儿教育), Canon(佳能)
(果叔) w/ knowledge Xianfeng Fruit(鲜丰水果), Fruitday(天天果园)
The Three Body similar movie w/o knowledge The Wandering Earth(流浪地球)
(三体) w/ knowledge Interstellar(星际穿越), Star Trek(星际迷航)
Gas Coupon product of prize w/o knowledge Fuel Card (加油卡)
(加油券) w/ knowledge Diesel(柴油), Gasoline(汽油), Gas Gift Card(加油礼品卡)
Tuxi Living Plus related company w/o knowledge Tuxi Catering(兔喜餐饮)
(兔喜生活+) w/ knowledge Carrefour(家乐福), CR Vanguard(华润), Walmart(沃尔玛)
4.3.2. Case Study I: How the Prior Knowledge Improves PAIR
In this section, we concentrate on the effectiveness of prior knowledge towards PAIR, which is expected to serve as strong evidence for relation filtering and entity expansion. We present the results of entity expansion for PAIR w/ knowledge and PAIR w/o knowledge in Table 3, respectively. Clearly, we find that target entities produced by the complete PAIR are more related to the corresponding source entities. On the contrary, PAIR w/o knowledge usually mine hallucinative entities (e.g., “Fruit Education” and “Tuxi Catering”) and fallacious entities (e.g., “Canon” with the entity “Uncle Fruit” and relation “related brand”), severely hindering the quality of MoKG.
4.3.3. Case Study II: How the Original Knowledge Graph is Enriched with PAIR
Since several efforts have been made for connecting users and various services / products in Alipay through hierarchical and spatio-related relations, PAIR could provide another complementary knowledge, i.e., marketing-related knowledge. Here, we take a close look at the novelty at the entity level by randomly selecting several specific entity types and presenting the average novelty in the original knowledge and MoKG with PAIR in Fig. 4. We find that the original knowledge graph is greatly enriched by PAIR, especially for the entity types “Book”, “Game” and “Disease”, which are rather rare before.
Figure 5. Deployment of LightPAIR (i.e., “Offline A”) for audience targeting and its comparison to the original EGL system with TRMP framework (Yang et al., 2023)(i.e., “Offline B”).
Table 4. Performance comparison for audience targeting with the metric # TAC (thousand). “RI” means the relative improvement of LightPAIR w.r.t. EGL.
Scenarios EGL LightPAIR RI
Uncle Fruit 7*1 8*7 +15.3%
The Three Body 3*3 3**6 +98.1%
Schwarzkopf 2*7 4*9 +93.1%
Biscuits voucher 1*2 1*3 +31.2%
Land Lords 9*2 2**2 +122.0%
Gas Coupon 3*8 6*1 +89.2%
4.3.4. Application: Audience Targeting
At last, we will demystify the practical application of MoKG in audience targeting, one of the fundamental issues of conducting marketing campaigns (Zhuang et al., 2020), which aims at effectively and efficiently locating target audiences who are interested in certain products / services / campaigns. As shown in Fig. 5, similar to the original EGL system with TRMP framework (Yang et al., 2023), the deployment of LightPAIR is decoupled with the online serving, thus satisfying the requirements of scalability and timeliness in industrial environments. We report the performance of LightPAIR and EGL based audience targeting in Table 4 with the metrics # TAC (i.e., the number of Target Audiences Covered). In Table 4, each scenario could be abstracted as several entities. Compared to the deployed EGL system, LightPAIR could achieve performance gains over the EGL by 15% - 122% in different marketing scenarios w.r.t. the number of Target Audiences Covered. Both of the observations show the effectiveness and efficiency of the proposed LightPAIR towards practical online marketing.
5. Conclusion
In this paper, we propose a novel PAIR and its lightweight version LightPAIR for harvesting marketing-oriented knowledge with LLMs, which are equipped with LLM based adaptive relation filtering and progressive prompting augmentation based entity expansion followed by a reliable aggregation that takes both self-consistency and semantic relatedness into consideration. Towards lightweight online serving, we specialize in a small and white-box (Light)PAIR, which is fine-tuned with high-quality corpora provided by a strong teacher-LLM. Extensive experiments demonstrate the effectiveness of the proposed (Light)PAIR via several key metrics for knowledge graph evaluation and practical applications in audience targeting also verify its applicability in real-world marketing scenarios. In the future, we would like to equip the proposed (Light)PAIR with the metapath-oriented entity expansion such that the expansion process can be more controllable and explainable.
Appendix A Examples of Progressive Prompting Augmentation
To further demonstrate the usage of different aspects of prior knowledge in the entity expansion procedure, we present a series of specific examples of progressive prompting augmentation in Figure 6 - 8, where different types of knowledge is distinguished by color, i.e., Struc_KG, Desc_KG and Inher_KG.
Figure 6. An example (source entity: Uncle Fruit, relation: related product) of calling an LLM to carry out progressive prompting augmentation.
Figure 7. An example (source entity: CKA, relation: related organization) of calling an LLM to carry out progressive prompting augmentation.
Figure 8. An example (source entity: air, relation: related brand) of calling an LLM to carry out progressive prompting augmentation.